You throw a ball straight up with a speed of 15 m/s. How long will it take the ball to come back to your hand?
1
Expert's answer
2012-10-10T09:16:27-0400
It will take

t = V0/g = 15[m/s]/9.8[m/s²] ≈ 1.53 seconds

to reach the top point and the same time will be elapsed to come back. So, the total time will be

T = 2t = 2*1.53[s] = 3.06 seconds.